A uniformly charged rod with total charge of -10 nC lies on the x-axis stretching from x = -3m to x = 3m. A point charge of magnitude Q = 15nc is positioned on the y-axis at the point (0m, 6m).

a. Find the electric field due to the charged rod at the point (8m,0). Your answer should be numeric (i.e. no symbols) and expressed in polar notation (i.e. magnitude and direction).

b. Find the electric field due to the point charge at the point (8m,0). Your answer should be numeric (i.e. no symbols) and expressed in polar notation (i.e. magnitude and direction).

c. Find the total electric field due to these charge distributions. i.e. find the vector sum of Part a. and Part b. Again, your answer should be expressed in polar form.

d. What is the electric potential energy at the same location (8m,0)?
Answer #1

- - - - - - 2 For such case E.F is given by. E TREÀ VI EcoBCP-a) to charge density our cale x-api A510710. 167x60c/ - Y-118 a1. (c) Net EF at point E = 33.75-3.75 | E= 30 Î N/ Electric pot = 8.p due to point charge TEP due to line charge v=E.E V = 33